OpenStudy (anonymous):

which exponential function goes through the points (1, 8) and (4, 521)?

OpenStudy (anonymous):

f(x) = 4(2)^x f(x) = 8(2)^-x f(x) = 4(4)^-x f(x) = 2(4)^x

OpenStudy (anonymous):

please help! ill give you medals

OpenStudy (mathmale):

You have at least two choices here: one would be to substitute the coordinates of the two given points into each of the given possible answers and determine which one of the answers is a true equation. The other approach would be to actually find the correct relationship yourself, using the model y=ax^b. Should you decide to do the problem in this manner, you'd need to find the values of a and b. Your choice.

OpenStudy (anonymous):

okay I got it, thank you

hero (hero):

I thought the general exponential formula was \(y = ab^x\)